# WENDY: Gene Regulatory Network Inference with Covariance Dynamics

### v1

@@ 5 Performance evaluation on experimental data
Besides synthetic data, there are also some GRNs determined by experiments, so that we can use them as the ground truth for the corresponding gene expression data [36, 68, 63]. In this section, we use two experimental data sets in Scenario 9 with known GRNs to test the performance of WENDY and three other GRN inference methods that work for different data types: GENIE3, NonlinearODEs, SINCERITIES. Since each cell is only measured once, we do not have the joint distribution of expression levels at different time points, and dynGENIE3 does not apply. WENDY ranks the first for the THP-1 data set, and ranks the third for the hESC data set.
@@ 5.1 THP-1 data set
THP-1 data set considers single-cell expression levels of monocytic THP-1 human myeloid leukemia cells, measured at $8$ time points, $t=$ 0, 1, 6, 12, 24, 48, 72, 96h, after applying phorbol-12-myristate-13-acetate [43]. For each time point, there are 120 cells measured. However, since each cell can be measured only once, we do not know how cells at different time points correspond to each other. Therefore, this data set is in Scenario 9.
The same as Papili Gao et al. [61], we consider the following 20 genes: BCL6, CEBPB, CEBPD, EGR2, FLI1, HOXA10, HOXA13, IRF8, MAFB, MYB, NFATC1, NFE2L1, PPARD, PPARG, PRDM1, RUNX1, SNAI3, TCFL5, TFPT, UHRF1. For such genes in THP-1 cells, there has been a GRN determined by experiments [75, 81] that we can use as the ground truth.
@@
(2) For SINCERITIES, we regard THP-1 data as Scenario 9 data. We use the data at $4$ time points: $t=$(1, 6, 12, 24).
(3) For NonlinearODEs, we regard THP-1 data as Scenario 8 data by taking average of gene expression levels over different cells. We use the data at $2$ time points: $t=$(0, 1).
(4) For GENIE3, we regard THP-1 data as Scenario 2 data by considering only one time point. We use the data at $t=48$.
We apply WENDY, SINCERITIES, NonlinearODEs, and GENIE3 to THP-1 data set and compare the corresponding AUROC and AUPR. See Table 5 for the results. We can see that WENDY is better than all other methods. Here GENIE3 has the worst performance, possibly because it cannot determine the sign (activation/inhibition) of a regulation. The results for SINCERITIES are slightly different from those reported by Papili Gao et al. [61], possibly because we do not consider diagonal elements of the GRN in the calculation of AUROC and AUPR.
@@
hESC data set considers single-cell expression levels of human embryonic stem cell-derived progenitor cells, measured at $6$ time points, $t=$0, 12, 24, 36, 72, 96h [19]. For each time point, there are 66–172 cells measured. Such cells are under development (which can be regarded as an intervention), and the regulation mechanism of their gene expression is highly complicated [46]. The same as THP-1 data set, there is no cell correspondence between different time points, and it is in Scenario 9.
Matsumoto et al. [53] calculated the mean expression level of each gene at each time point, and then calculated the variance of the mean expression levels at different time points for each gene. Using this approach, they selected out top 100 highly varying genes, which might actively regulate each other. Given the data of these 100 highly varying genes, we further select out genes that express in at least 95% of the cells. The reason is that when the expression data have too many 00s, graphical lasso might fail to converge, which fails WENDY. There are 18 genes selected: POU5F1, AEBP2, MIER1, SMAD2, ZNF652, ZFX, TERF1, SOX11, BBX, ZFP42, TULP4, ZNF471, ARID4B, ZNF483, SHOX, ZNF587, ZFP14, CEBPZ. For such genes, there has been a GRN determined by experiments [58, 73] that we can use as the ground truth.
Similar to the THP-1 data set, We apply WENDY (use $t=12,24$ data), SINCERITIES (use $t=0,12,24$ data), NonlinearODEs (use $t=24,36$ data), and GENIE3 (use $t=0$ data) to hESC data set and compare the corresponding AUROC and AUPR. See Table 6 for the results. We can see that WENDY is worse than SINCERITIES and NonlinearODEs.
@@ 6 Discussion
In this paper, we address the GRN inference problem for single-cell time series gene expression data following general interventions, where the joint distribution of different time points is unknown. Although this type of data is common in recent experiments, there are few GRN inference methods that fully utilize the information contained in the data. Therefore, we introduce WENDY, a GRN inference method developed for single-cell gene expression data spanning two time points after an intervention. This method is capable of inferring all mutual regulatory relations, including direction. We test WENDY and other GRN inference methods on two synthetic data sets and two experimental data sets, and the performance of WENDY is satisfactory. Besides, the time cost of WENDY is almost as low as the fastest method.
@@
The model in Subsection 3.1 only considers cells of the same type and under the same environment. In reality, cells of a complex organism are under different regulations by environmental factors, especially during development. For instance, during animal morphogenesis, retinoic acid can form a concentration gradient (positional information) [84], and leads cells at different positions to different fates by regulating certain genes [90]. If the levels of such spatially heterogeneous factors can be measured for each single cell, then one can add them into the GRN inference procedure. Otherwise, they become hidden variables in the gene expression model, and can make the inference much more difficult [50]. Besides such position-related factors, cell type can affect gene regulation [94]. One should label the cell type during experiments or after experiments using cluster analysis [9]. If the GRN inference is directly applied to data from different types of cells, the inference result might be unreliable [83]. In this work, we assume that cells are of the same type, and they are under the same environment.
@@
Although most gene expression models are similar to Eq. 1 that only considers the (mRNA) levels of $V_{1},\ldots,V_{n}$, some researchers argue that the actual gene expression mechanism should be more complex. In some models [10, 77, 35, 78], each gene can switch between “on” and “off” states, which correspond to different transcription rates. Some proteins can affect the transition rates between “on” and “off” states, which is the only way of gene mutual-regulation. When one gene is turned on, the number of its mRNA increases quickly, until it is turned off, when the mRNA count starts to decay exponentially. This leads to the mRNA bursts phenomenon [23], meaning that the mRNA count does not fluctuate locally, but has global bursts. Since proteins are translated from mRNAs, proteins also have bursts. However, proteins degrade much slower than mRNAs, meaning that mRNA count and protein count do not match exactly. Therefore, in such models, besides the mRNA counts $X_{1},\ldots,X_{n}$, there are also hidden variables for gene states and protein counts. However, more complicated models, generally using piecewise-deterministic Markov processes, are harder to solve, and researchers need to conduct large-scale simplifications. Besides, more complicated models need more accurate data, but the current mRNA measurement techniques are not very sensitive [97]. Therefore, we think that it is worthwhile to study differential equation based models.
@@
In our WENDY implementation, we employ the standard graphical lasso algorithm to determine the inverse of the covariance matrix. Other algorithms with similar purposes [80, 21] can also be considered for integration into WENDY.

### v2

@@ 5 Performance evaluation on experimental data
Besides synthetic data, there are also some GRNs determined by experiments, so that we can use them as the ground truth for the corresponding gene expression data (Hu et al.,, 2007; Reimand et al.,, 2010; Peter and Davidson,, 2015). In this section, we use two experimental data sets in Scenario 9 with known GRNs to test the performance of WENDY and three other GRN inference methods that work for different data types: GENIE3, NonlinearODEs, SINCERITIES. Since each cell is only measured once, we do not have the joint distribution of expression levels at different time points, and dynGENIE3 does not apply. WENDY ranks first for the THP-1 data set, and ranks second for the hESC data set.
@@ 5.1 THP-1 data set
THP-1 data set considers single-cell expression levels of monocytic THP-1 human myeloid leukemia cells, measured at $8$ time points, $t=$ 0, 1, 6, 12, 24, 48, 72, 96h, after applying phorbol-12-myristate-13-acetate (Kouno et al.,, 2013). For each time point, there are 120 cells measured. However, since each cell can be measured only once, we do not know how cells at different time points correspond to each other. Therefore, this data set is in Scenario 9.
The same as Papili Gao et al., (2018), we consider the following 20 genes: BCL6, CEBPB, CEBPD, EGR2, FLI1, HOXA10, HOXA13, IRF8, MAFB, MYB, NFATC1, NFE2L1, PPARD, PPARG, PRDM1, RUNX1, SNAI3, TCFL5, TFPT, UHRF1. For such genes in THP-1 cells, there has been a GRN determined by experiments (Tomaru et al.,, 2009; Vitezic et al.,, 2010) that we can use as the ground truth.
@@
(2) For SINCERITIES, we regard THP-1 data as Scenario 9 data. We use the data at $7$ time points: $t=$(1, 6, 12, 24, 48, 72, 96).
(3) For NonlinearODEs, we regard THP-1 data as Scenario 8 data by taking average of gene expression levels over different cells. We use the data at $3$ time points: $t=$(0, 1, 6).
(4) For GENIE3, we regard THP-1 data as Scenario 2 data by considering only one time point. We use the data at $t=24$.
We apply WENDY, SINCERITIES, NonlinearODEs, and GENIE3 to THP-1 data set and compare the corresponding AUROC and AUPR. See Table 5 for the results. We can see that WENDY is better than all other methods. Here GENIE3 has the worst performance, possibly because it cannot determine the sign (activation/inhibition) of a regulation. The results for SINCERITIES are slightly different from those reported by Papili Gao et al., (2018), possibly because we do not consider diagonal elements of the GRN in the calculation of AUROC and AUPR.
@@
hESC data set considers single-cell expression levels of human embryonic stem cell-derived progenitor cells, measured at $6$ time points, $t=$0, 12, 24, 36, 72, 96h (Chu et al.,, 2016). For each time point, there are 66–172 cells measured. Such cells are under development (which can be regarded as an intervention), and the regulation mechanism of their gene expression is highly complicated (Li and Belmonte,, 2017). The same as THP-1 data set, there is no cell correspondence between different time points, and it is in Scenario 9.
Matsumoto et al., (2017) calculated the mean expression level of each gene at each time point, and then calculated the variance of the mean expression levels at different time points for each gene. Using this approach, they selected out top 100 highly varying genes, which might actively regulate each other. Given the data of these 100 highly varying genes, we further select out genes that express in at least 95% of the cells. The reason is that when the expression data have too many 00s, graphical lasso might fail to converge, which fails WENDY. There are 18 genes selected: POU5F1, AEBP2, MIER1, SMAD2, ZNF652, ZFX, TERF1, SOX11, BBX, ZFP42, TULP4, ZNF471, ARID4B, ZNF483, SHOX, ZNF587, ZFP14, CEBPZ. For such genes, there has been a ground truth GRN determined by experiments (Neph et al.,, 2012; Stergachis et al.,, 2014) that we can use as the ground truth.
Similar to the THP-1 data set, we apply WENDY (use $t=12,24$ data), SINCERITIES (use $t=0,12,24,36,72$ data), NonlinearODEs (use $t=24,36,72,96$ data), and GENIE3 (use $t=24$ data) to hESC data set and compare the corresponding AUROC and AUPR. See Table 6 for the results. We can see that WENDY has the highest AUROC, but the total performance ranks second, worse than SINCERITIES.
@@ 6 Discussion
In this paper, we address the GRN inference problem for single-cell time series gene expression data following general interventions, where the joint distribution of different time points is unknown. Although this type of data is common in recent experiments, there are few GRN inference methods that fully utilize the information contained in the data. Therefore, we introduce WENDY, a GRN inference method developed for single-cell gene expression data spanning two time points after an intervention. This method is capable of inferring all mutual regulatory relations, including direction. We test WENDY and other GRN inference methods on two synthetic data sets and two experimental data sets, and the performance of WENDY is satisfactory: it is only weaker than dynGENIE3, which does not work on Scenario 9 data. See Table 7 for a summary. Besides, the time cost of WENDY is almost as low as the fastest method.
Table 7: Ranking of five GRN inference methods on four data sets
WENDY
SINCE-
RITIES
Nonline-
arODEs
GENIE3
dynG-
ENIE3
DREAM4 2nd 5th 4th 3rd 1st
SINC 3rd 5th 4th 1st 2nd
THP1 1st 2nd 3rd 4th -
hESC 2nd 1st 4th 3rd -
@@
The model in Subsection 3.1 only considers cells of the same type and under the same environment. In reality, cells of a complex organism are under different regulations by environmental factors, especially during development. For instance, during animal morphogenesis, retinoic acid can form a concentration gradient (positional information) (Wang et al.,, 2020), and leads cells at different positions to different fates by regulating certain genes (Wolpert,, 1989). If the levels of such spatially heterogeneous factors can be measured for each single cell, then one can add them into the GRN inference procedure. Otherwise, they become hidden variables in the gene expression model, and can make the inference much more difficult (Lo et al.,, 2015). Besides such position-related factors, cell type can affect gene regulation (Zhang et al.,, 2023). One should label the cell type during experiments or after experiments using cluster analysis (Bocci et al.,, 2022). If the GRN inference is directly applied to data from different types of cells, the inference result might be unreliable (Wang and He,, 2023). In this work, we assume that cells are of the same type, and they are under the same environment. An important future direction is to build more realistic gene expression models and develop corresponding GRN inference methods.
@@
Although most gene expression models are similar to Eq. 1 that only considers the (mRNA) levels of $V_{1},\ldots,V_{n}$, some researchers argue that the actual gene expression mechanism should be more complex. In some models (Bonnaffoux et al.,, 2019; Ventre,, 2021; Herbach,, 2023; Ventre et al.,, 2023), each gene can switch between “on” and “off” states, which correspond to different transcription rates. Some proteins can affect the transition rates between “on” and “off” states, which is the only way of gene mutual-regulation. When one gene is turned on, the number of its mRNA increases quickly, until it is turned off, when the mRNA count starts to decay exponentially. This leads to the mRNA bursts phenomenon (Elgart et al.,, 2011), meaning that the mRNA count does not fluctuate locally, but has global bursts. Since proteins are translated from mRNAs, proteins also have bursts. However, proteins degrade much slower than mRNAs, meaning that mRNA count and protein count do not match exactly. Therefore, in such models, besides the mRNA counts $X_{1},\ldots,X_{n}$, there are also hidden variables for gene states and protein counts. However, more complicated models, generally using piecewise-deterministic Markov processes, are harder to solve, and researchers need to conduct large-scale simplifications. Besides, more complicated models need more accurate data, but the current mRNA measurement techniques are not very sensitive (Zhang et al., 2021b, ). Therefore, we think that it is worthwhile to study differential equation based models.
@@
In our WENDY implementation, we employ the standard graphical lasso algorithm to determine the inverse of the covariance matrix. Other algorithms with similar purposes (Vinci et al.,, 2019; Danaher et al.,, 2014) can also be considered for integration into WENDY.